# A Deep Learning Approach for Pixel-level Material Classification via Hyperspectral Imaging

### v1

@@ 1 Introduction
To this end, Section 2 describes the related work, while Section 3 provides a high-level description of the system setup and the hardware equipment used in this work. The creation of hyperpsectral dataset, the spectral preprocessing methodologies along with the semi-automated AI-assisted mask generation are presented in Section 4. Section 5 introduces the architecture of the proposed model, the preparation of the training instances. The training process process, along with the selected hyperparameters and the initial results are also discussed in that section. In addition, the experimental results on various scenarios, i.e. mixed materials, shredded samples and overlapping objects, and the performance of the proposed system are discussed in Section 6. This section, also, explores the limitations of the hyperspectral model in the case of dark-coloured or black samples. Finally, in Section 7 a discussion about the main conclusions of this study and the future steps are provided.
2 Related Work
@@
The combination of deep learning and hyperspectral imaging has shown great promise in improving the material classification capability in Earth Observation applications. Notable examples include models, e.g. Xception-based, CNN-based systems, and R-VCANet achieving accuracies up to 99%, with precision and recall values around 94% and 93%, and F1-scores reaching 91% [23, 24, 25, 26, 27]. Venkatesan et al. [28] applied deep recurrent neural networks in medical hyperspectral images for feature recognition achieving a precision rate of about 96%. Xiong et al. [29] developed material tracking methods for hyperspectral videos using deep learning, achieving an F1-score of 94%. Medus et al. [30] applied CNNs to classify hyperspectral images in industrial food packaging, reporting an accuracy of 99%. Okada et al. followed a patch-wise approach for the identification of 5 different mineral types, using a VGG16-based CNN to classify the acquired HS images they achieved high accuracy over 90% [31]. Extreme Learning combined with Stacked Autoencoders for feature extraction has also been applied on HS images, in [32], for the detection of plastic films within cotton feed stock, creating a pixel-level classification map with accuracy up to 95%. Zhu et al. in [33], also attempted to classify different cotton seeds. Their work encompasses a multivariate analysis for manual extraction of the 10 most informative features, which were subsequently fed to CNN for the classification task, resulting in a prediction accuracy of 88%. Finally, Artzai et al. worked with HS images with 76 bands of non ferrous metals, and utilised them to train a CNN-based U-Net like network [34]. This approach analysed the HS images as a whole and created a classification map with an accuracy equal to 95%.
@@
The aforementioned studies with hyperspectral classification using deep learning have yielded remarkable results. However, it is important to note most of the already limited approaches that exist on the field commonly either deal with HS images in batches or utilise 2D Convolutional Neural Networks (CNNs). In order to achieve this, they use 2D convolutional kernels that inevitably lead to loss of accuracy at object borders. Notably, for 2D CNNs convolutions are designed to capture spatial context within each individual frame but this may blur boundaries or misclassify pixels on the boundary between neighboring objects that have different spectral signatures. This means that spectral continuity across edges is not incorporated by the 2D convolution operation causing edge artifacts and reducing classification accuracy here. These limitations highlight the need for more refined techniques, that analyse HS images on a pixel level and can preserve boundary integrity while leveraging the rich spectral data provided by hyperspectral imaging.
3 Hyperspectral Imaging Setup
@@ 4.1 Dataset description
Train Set: In detail, 169 objects were selected to represent the classes. The physical objects were separated according to their material classes and subsequently placed randomly on the conveyor belt, described in Section 3.2. For the augmentation of the training set by every sample was captured twice with slightly different light conditions. A comprehensive analysis of the dataset composition is presented in Table I.
@@
TABLE III: Black samples test subset.
@@ 4.2 Acquisition Pipeline
It should be noted that in HS image processing applications it is highly recommended to use the aforementioned steps for spectral calibration and normalization which often include expensive equipment and dependencies on the hardware manufacturer. However, as shown in Section 6, even if we totally ignore those steps, the learning performance of the proposed deep learning algorithm is not affected.
@@
4.3 Ground truth mask generation
@@
5 Model Architecture & Training
5.1 Architecture
@@ 6 Results & Discussion
As discussed in Section 4 the test-set presented in Table III includes images of all the HDPE, PET, PP, PS classes, which however have not been utilised in the training subset. In this manner, guarantee is provided that no same object or pixel is simultaneously evident in both train and test data.
@@
To quantify the model’s performance, the accuracy score was calculated for the whole test set in pixel level thus providing a detailed evaluation on the model’s classification capacity. The overall accuracy achieved by the model is 97.44%. The confusion matrix is presented in Figure 8, summarizing the classification performance among different material classes.
@@
The cell values in the confusion matrix are row-wise normalized, i.e. normalized with respect to the total number of samples in each class. By observing the confusion matrix of Figure 8 is easy to notice that less than 1% of the total pixels of each class is falsely classified as another material and most of the errors are misclassifications between materials and the Background.
It can be calculated that 97.45% of the error comes in the borders between objects and the background. After more careful inspection is not easy to conclude that the error is actually on the predictions of P1CH and not on the ground truth masks. Figure 9 presents two zoomed-in crops of Figure 7, where it is clearly shown that in both cases the P1CH has the capacity to generate much smoother and precise masks for the respective object while simultaneously predicting correctly the material class. Comparing, also, the zoomed crops one can realize the mistake made on the top right side of the PP sample’s mask, as well as at the PET bottle’s spout. In both cases, P1CH was capable of predicting a more precise than the semi-automatically generated mask.
@@
7 Conclusion & Future Work
In this paper a lightweight 1D Convolutional Neural Hyperspectral Classification, system, P1CH, is proposed. The developed model generates highly detailed and precise classification maps. Unlike conventional RGB methods, the proposed algorithm utilises the spectral information encoded in hyperspectral image’s pixel, allowing to detect and correctly classify objects in challenging scenarios.
@@ 7 Conclusion & Future Work
Acknowledgements
This research was financially supported by the European Union’s Horizon Europe research and innovation program under grant agreement No 101138789 (project W2W).

### v2

@@ I Introduction
To this end, Section II describes the related work, while Section III provides a high-level description of the system setup and the hardware equipment used in this work. The creation of hyperpsectral dataset, the spectral preprocessing methodologies along with the semi-automated AI-assisted mask generation are presented in Section IV. Section V introduces the architecture of the proposed model, the preparation of the training instances. The training process process, along with the selected hyperparameters and the initial results are also discussed in that section. In addition, the experimental results on various scenarios, i.e. mixed materials, shredded samples and overlapping objects, and the performance of the proposed system are discussed in Section VI. This section, also, explores the limitations of the hyperspectral model in the case of dark-coloured or black samples. Finally, in Section VII a discussion about the main conclusions of this study and the future steps are provided.
II Related Work
@@
The combination of deep learning and hyperspectral imaging has shown great promise in improving the material classification capability in Earth Observation applications. Notable examples include models, e.g. Xception-based, CNN-based systems, and R-VCANet achieving accuracies up to 99%, with precision and recall values around 94% and 93%, and F1 score reaching 91% [23, 24, 25, 26, 27]. Venkatesan et al. [28] applied deep recurrent neural networks in medical hyperspectral images to achieve feature recognition with a precision rate of about 96%. Xiong et al. [29] developed material tracking methods for hyperspectral videos using deep learning, achieving an F1 score of 94%. Medus et al. [30] applied CNNs to classify hyperspectral images in industrial food packaging, reporting an accuracy of 99%. Okada et al. followed a patch-wise approach for the identification of 5 different mineral types, using a VGG16-based CNN to classify the acquired HS images they achieved high accuracy over 90% [31]. Extreme Learning combined with Stacked Autoencoders for feature extraction has also been applied on HS images, in [32], for the detection of plastic films within cotton feed stock, creating a pixel-level classification map with accuracy up to 95%. Zhu et al. in [33], also attempted to classify different cotton seeds. Their work encompasses a multivariate analysis for manual extraction of the 10 most informative features, which were subsequently fed to CNN for the classification task, resulting in a prediction accuracy of 88%. Moreover, Artzai et al. worked with HS images with 76 bands of non ferrous metals, and utilised them to train a CNN-based U-Net like network [34]. This approach analysed the HS images as a whole and created a classification map with an accuracy equal to 95%. Finally, Roy et al.[35] in their work proposed a methodology that incorporates PCA as a preprocessing step with a deep convolutional network, i.e. HybridSN, for the pixel-wise classification of Hyperspectral Images, achieving an accuracy score of 99%.
@@
The aforementioned studies with hyperspectral classification using deep learning have yielded remarkable results. However, it is important to note that most of the already limited approaches that exist in the field commonly either deal with HS images in batches or utilise 2D Convolutional Neural Networks (CNNs). In order to achieve this, they use 2D convolutional kernels that inevitably lead to loss of accuracy at object borders. Notably, for 2D CNNs convolutions are designed to capture spatial context within each individual frame but this may blur boundaries or misclassify pixels on the boundary between neighboring objects that have different spectral signatures. This means that spectral continuity across edges is not incorporated by the 2D convolution operation, causing edge artifacts and reducing classification accuracy here. These limitations highlight the need for more refined techniques that analyse HS images on a pixel level and can preserve boundary integrity while leveraging the rich spectral data provided by hyperspectral imaging.
III Hyperspectral Imaging Setup
@@ IV-A Dataset description
Train Set: In detail, 169 objects were selected to represent the classes. The physical objects were separated according to their material classes and subsequently placed randomly on the conveyor belt, described in Section III-B. For the augmentation of the training set by every sample was captured twice with slightly different light conditions. A comprehensive analysis of the dataset composition is presented in Table I.
@@
TABLE III: Black samples test
subset.
@@ IV-B Acquisition Pipeline
It should be noted that in HS image processing applications it is highly recommended to use the aforementioned steps for spectral calibration and normalization which often include expensive equipment and dependencies on the hardware manufacturer. However, as shown in Section VI, even if we totally ignore those steps, the learning performance of the proposed deep learning algorithm is not affected.
@@
IV-C Ground truth mask generation
@@
V Model Architecture & Training
V-A Architecture
@@ VI Results & Discussion
As discussed in Section IV the test-set presented in Table III includes images of all the HDPE, PET, PP, PS classes, which however have not been utilised in the training subset. In this manner, guarantee is provided that no same object or pixel is simultaneously evident in both train and test data.
@@
To quantify the model’s performance, the accuracy score was calculated for the whole test set in pixel level thus providing a detailed evaluation on the model’s classification capacity. The overall accuracy achieved by the model is 97.44%. The confusion matrix is presented in Figure 8, summarizing the classification performance among different material classes. The cell values in the confusion matrix are row-wise normalized, i.e. normalized with respect to the total number of samples in each class. By observing the confusion matrix of Figure 8 is easy to notice that less than 1% of the total pixels of each class is falsely classified as another material and most of the errors are misclassifications between materials and the Background.
@@
To further validate the performance of the proposed P1CH model, the state-of-the-art HybridSN methodology [35, 49] was replicated and evaluated on the same test dataset. This approach ensures a direct and fair comparison under identical conditions. The overall accuracy achieved by the HybridSN model on this dataset is significantly lower, at 21.81%. In contrast, the P1CH model achieves an accuracy of 97.44%, along with superior recall and Kappa score. These results, summarized in Table IV, underline the robustness and reliability of the proposed P1CH model in pixel-level material classification tasks.
@@
TABLE IV: Performance Metrics Comparison between P1CH and HybridSN Models
Metric P1CH HybridSN
Mean Accuracy (%) 97.44 21.81
Mean Recall (%) 85.99 7.86
Mean Kappa Score 0.9295 -0.0795
Mean Inference Time (sec) 5.06 34.29
The performance metrics summarized in Table IV further underscore the effectiveness of the proposed P1CH classifier. Notably, the accuracy achieved by P1CH (97.44%) is significantly higher than that of the HybridSN model (21.81%), highlighting its superior capacity to correctly classify pixels across the test set. In addition to accuracy, the recall value of P1CH (85.99%) demonstrates its robustness in identifying material classes with high consistency, in stark contrast to the HybridSN model, which achieved a recall of only 7.86%. The Kappa score further illustrates the reliability of the P1CH model, with a value of 0.9295 compared to the negative score produced by HybridSN, reflecting its struggles with pixel-level classification tasks.
Beyond classification accuracy, the computational efficiency of P1CH is a notable advantage. The proposed model achieves an inference speed of 200 lines per second (5.06 seconds per image), far surpassing HybridSN’s 29 lines per second (34.29 seconds per image). This substantial difference not only demonstrates the lightweight design of P1CH but also makes it a practical choice for almost real-time applications, where rapid and reliable pixel-level classification is essential. With these results, P1CH emerges as a robust, accurate, and highly efficient solution for hyperspectral image pixel-wise classification in demanding environments.
A visual comparison of predictions by both models, on the test set, is presented in Figure 10. The HybridSN model struggles with pixel-level classification accuracy, particularly at object boundaries, and often misclassifies a large portion of pixels. In contrast, the P1CH model demonstrates precise segmentation and classification, even in challenging scenarios involving overlapping objects or small irregular shapes. These findings validate the effectiveness of the P1CH model and its potential applicability in real-world hyperspectral imaging tasks.
Figure 10: From left to right: The false-colour HS image, the Ground Truth mask, the classification maps generated by HybridSN model ,as well as the generated classification map, by the P1CH Classifier, in the aforementioned scenarios.
Focusing on the borders
Moreover, it can be calculated that 97.45% of the error comes in the borders between objects and the background, as suggested by the relatively low recall score. After more careful inspection is not easy to conclude that the error is actually on the predictions of P1CH and not on the ground truth masks. Figure 9 presents two zoomed-in crops of Figure 7, where it is clearly shown that in both cases the P1CH has the capacity to generate much smoother and precise masks for the respective object while simultaneously predicting correctly the material class. Comparing, also, the zoomed crops one can realize the mistake made on the top right side of the PP sample’s mask, as well as at the PET bottle’s spout. In both cases, P1CH was capable of predicting a more precise than the semi-automatically generated mask.
@@
VII Conclusion & Future Work
In this paper a lightweight 1D Convolutional Hyperspectral Classification, system, P1CH, is proposed. The developed model generates highly detailed and precise classification maps. Unlike conventional RGB methods, the proposed algorithm utilises the spectral information encoded in hyperspectral image’s pixel, allowing to detect and correctly classify objects in challenging scenarios.